OpenStudy (anonymous):

A statue at the zoo casts a shadow of 6 feet. On the other side of the zoo, an 11-foot elephant casts a 13-foot shadow. How tall is the statue?

OpenStudy (anonymous):

set this up like a proportion: \[\frac{ 6 }{ x }=\frac{ 13 }{ 11 }\] Simply cross multiply .. so 6 * 11 = ? then divide that number by 13, and you'll have your answer